OpenStudy (anonymous):

Need help final test What is the solution to the equation 4(2b-4)=2(3-5b) solve for b

jigglypuff314 (jigglypuff314):

distributive property is like a(b-c) = ab - ac so use distributive property for that... 4(2b-4) = 2(3-5b) -> 8b - 16 = 6 - 10b then you can add 10b to both sides then add 16 to both sides then divide both sides by 18 and reduce/simplify :)

OpenStudy (anonymous):

Thanks your a life saver

jigglypuff314 (jigglypuff314):

you're welcome :)